OpenStudy (magbak):

Find the perimeter of a rectangular object which has a length of square root of 128 feet and a width of square root of 200 feet.

OpenStudy (jhannybean):



OpenStudy (jhannybean):

the formula for Perimeter follows \[\large P=2l+2w\] Where \(l = \text{length} \ \text{and} \ w=\text{width}\)

OpenStudy (jhannybean):

in our problem, \(\large l =\sqrt{128} \ , \ w =\sqrt{200}\)

OpenStudy (jhannybean):

Understand how it works? :)

OpenStudy (jhannybean):

Lets see. \[\large P = 2(\sqrt{128})+2(\sqrt{200})\]\[\large P= 2\sqrt{128} + 2\sqrt{200}\]

OpenStudy (jhannybean):

Is your answer in decimal form?

OpenStudy (magbak):

No decimal form is not accepted

OpenStudy (jhannybean):

Ok.

OpenStudy (magbak):

No there is no answer choices it is a written response do you want the entire question? ;)

OpenStudy (magbak):

Part 1: Find the perimeter of a rectangular object which has a length of square root of 128 feet and a width of square root of 200 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (jhannybean):

Oh nevermind. Well,we have to factor out 128 and 200, once we make the stuff inside the square root the same base, we can combine them :)

OpenStudy (jhannybean):

so \[\large P= 2\sqrt{128} + 2\sqrt{200}= 2\sqrt{2\cdot 2^6}+2\sqrt{2^3\cdot5^2}\]

OpenStudy (jhannybean):

Do you know how to pull out pairs of numbers from under the parenthesis?

OpenStudy (magbak):

Their is no parenthesis

OpenStudy (magbak):

OK thank you I will solve the rest form here. I appreciate your concern. Thank you very much. :)

OpenStudy (jhannybean):

not parenthesis,imean square root*

OpenStudy (jhannybean):

\[\large 2\cdot (2^{1/2}\cdot 2^{6/2})+2\cdot (2^{3/2}\cdot 5^{2/2})\]You see what powers can be easily reduced and pull them out of the parenthesis. \[\large 2\cdot 2^3 (2^{1/2}) + 2\cdot 5 (2^{3/2})\] rewrite \[\large 16\sqrt{2} +10\sqrt{2^3}\] and nw we see that we can pull out one more 2 from 2^3 since 2^3 =2 x 2 x 2 \[\large 16\sqrt{2} +10\sqrt{2\cdot 2\cdot 2}= 16\sqrt{2}+10\cdot 2\sqrt{2} \]Simplify \[\large 16\sqrt{2}+20\sqrt{2} \] the stuff under the square root is the same,or what we call "the base" so we can add these two together. \[\large 16+20\sqrt{2} =?\]

OpenStudy (magbak):

Ok thank you I got the answer thanks .

OpenStudy (jhannybean):

Awesome :D